# iTBLS: A Dataset of Interactive Conversations Over Tabular Information

### v1

@@ Abstract
This paper introduces Interactive Tables (iTBLS), a dataset of interactive conversations situated in tables from scientific articles. This dataset is designed to facilitate human-AI collaborative problem-solving through AI-powered multi-task tabular capabilities. In contrast to prior work that models interactions as factoid QA or procedure synthesis, iTBLS broadens the scope of interactions to include mathematical reasoning, natural language manipulation, and expansion of existing tables from natural language conversation by delineating interactions into one of three tasks: interpretation, modification, or generation. Additionally, the paper presents a suite of baseline approaches to iTBLS, utilizing zero-shot prompting and parameter-efficient fine-tuning for different computing situations. We also introduce a novel multi-step approach and show how it can be leveraged in conjunction with parameter-efficient fine-tuning to achieve the state-of-the-art on iTBLS; outperforming standard parameter-efficient fine-tuning by up to 15% on interpretation, 18% on modification, and 38% on generation.
@@ 1 Introduction
Recent research on Conversational AI has focused on adding enhanced multi-task capabilities to large language models (LLMs). This research includes building systems capable of situated interactions over structured knowledge sources such as tabular information Sundar and Heck (2022). Automated methods for tabular interpretation, manipulation, and generation empower users by saving time and reducing errors in managing tabular content Kardas et al. (2020). Previous studies have focused on individual aspects of tabular data management: representation learning for interpretation tasks like grounded question answering, manipulation for data wrangling, and generation for summarizing textual information independently Nakamura et al. (2022a); Sundar and Heck (2023); Sundar et al. (2024); Fang et al. (2024).
@@
To address these challenges, this paper introduces Interactive Tables (iTBLS)  [^1] The dataset and baseline models will be released with the camera-ready version under Apache License 2.0, a dataset of interactive conversations in English situated in tabular information. iTBLS decomposes the challenge into three distinct tasks: interpretation, which involves understanding tabular content within a conversational framework; modification, which entails manipulating tabular content through natural language commands; and generation, which focuses on integrating new natural language information into existing tables. The tabular information in iTBLS is sourced from scientific articles hosted on arXiv  [^2] https://arxiv.org, an open-access repository of academic preprints.
In contrast to earlier methods that approached tabular interpretation as question-answering focused on single-cell factoids, iTBLS introduces a more complex conversational interaction framework. Beyond factoid question-answering, iTBLS encompasses tasks such as comparison, determining absolute and relative positions, and mathematical reasoning. Previous research primarily examined procedural command generation for spreadsheets or the alignment of tabular data through LLMs. iTBLS integrates these functionalities into a unified task, enabling the manipulation of existing tables through natural-language commands. On tabular generation, while prior work addressed the summarization of natural language paragraphs in a tabular format, iTBLS focuses on generating row or column data conversationally.
Furthermore, this paper presents baseline approaches utilizing zero-shot prompting of off-the-shelf LLMs and parameter-efficient fine-tuning and represents a suite of techniques for different computing situations. Additionally, a novel multi-step approach is presented to improve accuracy and reduce errors, inspired by past work using multi-step methods to improve a variety of approaches including complex reasoning Jiang et al. (2023b), external tool use Paranjape et al. (2023), feedback Richardson et al. (2023), self-refinement Madaan et al. (2024), and mathematical reasoning Zhou et al. (2023). This multi-step method results in an improvement over standard parameter-efficient fine-tuning by up to 15% on interpretation, 18% on modification, and 38% on generation.
@@
  • Creating iTBLS, a dataset of tabular interactions unifying interpretation, modification, and generation
@@
  • Broadening the scope of interactions to include mathematical reasoning, natural language manipulation, and natural language expansion
  • Completing a series of baselines utilizing zero-shot prompting, parameter-efficient fine-tuning, and a task-specific multi-step approach
The rest of the paper is organized as follows: Section 2 reviews prior research addressing tabular interpretation, modification, and generation. Section 3 presents a description of the tasks in iTBLS and outlines the dataset collection process. Sections 4 and 5 detail the baseline approaches to address iTBLS and their associated results. Lastly, Section 6 concludes the paper.
@@ 2.1 Tabular Interpretation
Tabular interpretation involves a dialogue turn focused on extracting information from a specific cell in a table, such as identifying a cell satisfying certain criteria. Prior research on tabular interpretation focused on grounded question-answering. An important challenge in the collection of such datasets is the availability of large-scale tabular data. Consequently, many tabular datasets are constructed from online resources such as Wikipedia including WikiTableQuestions Pasupat and Liang (2015), ManyModalQA Hannan et al. (2020), TaBERT Yin et al. (2020), NQ-Tables Herzig et al. (2021), FEVEROUS Aly et al. (2021), FeTaQA Nan et al. (2022), HybriDialogue Nakamura et al. (2022b), and HiTab Cheng et al. (2022). Other tabular datasets are constructed from financial reports including TAT-QA Zhu et al. (2021), FinQA Chen et al. (2021), MultiHiertt Zhao et al. (2022).
Proposed approaches to address the tabular interpretation task include architectures based off of the Transformer encoder Yin et al. (2020); Herzig et al. (2020); Chen et al. (2020); Eisenschlos et al. (2020); Liu et al. (2021); Gu et al. (2022); Yang et al. (2022), decoder Gong et al. (2020); Akhtar et al. (2023); Zha et al. (2023); Jiang et al. (2023a); Zhang et al. (2023b); Sui et al. (2024), or both (encoder-decoder) Nakamura et al. (2022b); Deng et al. (2022); Sundar and Heck (2023); Sarkar and Lausen (2023).
@@ 2.2 Tabular Modification
Tabular modification concerns the manipulation of the content within an existing table without altering the overall structure of rows and columns. Early work on tabular modification explored the generation of procedural commands for spreadsheets using synthesis algorithms Singh and Gulwani (2012); Shigarov and Mikhailov (2017). Tools utilizing programming-by-example to parse user intents into executable commands have also been explored Scaffidi et al. (2009); Kandel et al. (2011); Jin et al. (2017); Petricek et al. (2023); Chen et al. (2023). More recent work has shifted focus towards leveraging LLMs to synthesize commands for tools Huang et al. (2023), reformat tabular information Nobari and Rafiei (2023), and execute programming commands Liu et al. (2023).
@@ 2.3 Tabular Generation
Tabular generation focuses on expanding an existing table by adding a new row or column. Research on tabular generation initially employed discriminative techniques, such as tree-based methods for generating tables of contents Branavan et al. (2007) and SVMs to classify text across various labels Aramaki et al. (2009). Recent approaches have shifted towards neural techniques including Generative Adversarial Networks (GANs) Xu and Veeramachaneni (2018); Park et al. (2018); Chen et al. (2019); Zhao et al. (2021), Autoencoders Li et al. (2019); Darabi and Elor (2021), Diffusion models Kotelnikov et al. (2023), and LLMs Borisov et al. (2023); Solatorio and Dupriez (2023); Zhang et al. (2023a); Gulati and Roysdon (2023); Zhao et al. (2023); Seedat et al. (2024).
A similar line of research also explores the generation of tabular data from associated textual information. Wu et al. (2022) introduced four datasets and proposed a modification to the Transformer’s attention mechanism to summarize textual information in a tabular format by inverting datasets created for the dual task of converting tables to text, (as opposed to new conversational evidence). Other approaches to summarize textual information in a tabular format include the addition of learnable bias parameters Pietruszka et al. (2022), structure-aware instruction-tuning Tang et al. (2023), and reformulating the task within a question-answering framework Sundar et al. (2024).
@@ 3 The iTBLS Dataset
The Interactive Tables (iTBLS) dataset features 3-turn conversational interactions situated in tabular data, covering the three distinct types of interactions described in Section2: interpretation, modification, and generation. Each example type is exemplified in Figure 1 and described below. In addition, since the mode of interaction is not known a priori, any proposed approach using iTBLS must effectively identify the interaction type, either explicitly or implicitly. In the following sections, we provide a detailed description of each type of interaction and outline the dataset collection process.
@@ 3.1 Tasks
Tabular Interpretation: In iTBLS, interpretive interactions are structured as question-answer pairs, where the goal is to identify the cell referred to by the question. The references could be absolute (referring to a specific row or column), or relative (referring to one cell in the context of another). Appendix A.1 details absolute and relative references in iTBLS.
Tabular modification: We conceptualize modification in iTBLS as a series of cell swaps, positing that any content rearrangement can ultimately be reduced to such exchanges. This approach allows for both explicit references, where specific row and column numbers are cited, and implicit references, which rely on the content or relative positions of cells. Table 7 in Appendix A.1 showcases examples from iTBLS. As observed, there is a mix of explicit and implicit references to the specific contents to be manipulated.
Tabular generation: In iTBLS, table generation is guided by new natural language evidence. This evidence clarifies appending a row or column, defines the suitable header, and supplies the data entries for the new row (or column) relative to existing columns (or rows). This process ensures that the added elements are contextually relevant and accurately integrated into the table. Table 8 in Appendix A.1 provides examples of such interactions, demonstrating how users can request the incorporation of new row and column data into an established table framework.
@@ 3.2 Dataset Collection
For our dataset creation, we employ two distinct sources for generating dialogue turns based on the type of interaction and the specific table component involved. For tasks related to tabular interpretation and modification, we engage crowd-workers from Amazon Mechanical Turk (AMT). These workers are tasked with formulating questions or commands that pertain to the pre-identified cell(s) designated for each interaction. We recruit workers from Australia, Canada, Ireland, New Zealand, the United Kingdom, and the USA. Each crowdworker is compensated at a rate of \$0.15 per Human Intelligence Task (HIT), with the average completion time for each HIT being approximately 40 seconds. Detailed information on the AMT interface used for these tasks is included in Appendix A.3.
For generation, GPT-4 is prompted to write a dialogue turn summarizing a row or column of the table. We prompt the version of GPT-4 as of 2/16/2024 using the OpenAI API. The prompt is as follows:
@@
Each sample in the dataset contains the source arXiv ID, the table that the conversation is situated in, the index of that table within the paper (e.g. Table X), the utterance describing the interaction, the ground truth cell(s) involved in the interaction, and finally the expected output. Statistics of the dataset are provided in Table 1.
Interpret Modify Generate
train dev test train dev test train dev test
# Samples 3001 646 521 3001 646 521 3001 646 521
@@
Tokens 14.3 14.3 14.3 18.5 17.7 17.8 59.3 60.6 55.9
Words 10.7 10.7 10.3 13.5 13.2 13.0 32.0 31.8 29.5
# Per Input Table
Cells 28.0 28.8 27.7 28.0 28.8 27.5 25.2 26.0 25.1
Columns 5.0 5.0 4.9 5.0 5.0 4.9 4.8 4.8 4.7
Rows 5.5 5.5 5.6 5.5 5.5 5.6 5.3 5.3 5.3
Table 1: Dataset Statistics
Figure 2: iTBLS multi-step approach. The input table and user request are first converted into an encoded form by an LLM. The encoded request is then used to generate the output via a task-specific procedure in python.
4 Baseline Approaches
[…]
4.1 Zero-shot prompting
Zero-shot prompting leverages off-the-shelf LLMs that have undergone pre-training with a causal language modeling objective followed by supervised fine-tuning to align with human instructions. The LLM is not fine-tuned on iTBLS and is directly prompted without examples in-context. The prompt consists of the input table, the interaction request, and additional task-specific instructions to ensure the output is formatted correctly.
4.2 Parameter-efficient fine-tuning (PEFT)
We use QLoRA Dettmers et al. (2023), a parameter-efficient approach to fine-tune LLMs. In this method, an LLM is fine-tuned for each sub-task by maximizing a causal language modeling objective based on the specific interaction. The objective to be optimized is described in Section 5.
4.3 Multi-step approach
The Multi-Step approach (MS) breaks down the entire task into a series of intermediate steps, addressing part of the specific task. For interpretation and modification, this process first identifies the cells referred to by the user using an LLM in a zero-shot configuration followed by program synthesis to execute the action. However, due to the custom encoding used in identifying cell references (Figure 2), the approach did not yield satisfactory results when compared to zero-shot prompting. For generation, the multi-step approach breaks down table generation as two-step question-answering. Additional details are provided in Section 5.3.
4.4 Multi-Step + PEFT
This approach first breaks down the entire process into a sequence of intermediate steps. Next, an LLM is fine-tuned using parameter-efficient fine-tuning to address each step independently.

### v2

@@ Abstract
This paper introduces Interactive Tables (iTBLS), a dataset of interactive conversations that focuses on natural-language manipulation of tabular information sourced from academic pre-prints on ArXiv. The iTBLS dataset consists of three types of tabular tasks – interpretation, modification, and generation. Interpretation focuses on tabular understanding, modification focuses on manipulating tabular information, and generation focuses on the addition of new natural-language evidence. In addition, the paper presents a novel framework that reformulates tabular operations as question-answering, where an appropriate question is formulated based on the nature of interaction and the question is answered using the user request as evidence. The developed approach results in an improvement on all tasks on a sequence-to-sequence modeling baseline on iTBLS. In addition, the question-answering-based reformulation is applied to datasets from prior work for the text-to-table task where textual paragraphs are summarized into tables. The novel approach results in up to 13% improvement in Exact-Match accuracy and up to 16% improvement in BERTScores compared to the prior state-of-the-art.
iTBLS: A Dataset of Interactive Conversations Over Tabular Information
Anirudh Sundar1  and Christopher Richardson2 thanks: Work done while at Georgia Tech  and Adar Avsian1  and Larry Heck1 1 Georgia Institute of Technology, USA 2 Google Inc., USA asundar34, larryheck@gatech.edu
@@ 1 Introduction
Recent research on Conversational AI has focused on adding enhanced multi-task capabilities to large language models (LLMs). This research includes building systems capable of situated interactions over structured knowledge sources such as tabular information Sundar and Heck (2022). Automated methods for tabular interpretation, manipulation, and generation empower users by saving time and reducing errors in managing tabular content Kardas et al. (2020). Previous studies have focused on individual aspects of tabular data management: representation learning for interpretation tasks like grounded question answering, manipulation for data wrangling, and generation for summarizing textual information independently Nakamura et al. (2022a); Sundar and Heck (2023); Fang et al. (2024).
@@
To address these challenges, this paper introduces Interactive Tables (iTBLS)  [^1] https://huggingface.co/datasets/avalab/iTBLS, a dataset of interactive conversations in English situated in tabular information. iTBLS decomposes the challenge into three distinct tasks: interpretation, which involves understanding tabular content within a conversational framework; modification, which entails manipulating tabular content through natural language commands; and generation, which focuses on integrating new natural language information into existing tables. The tabular information in iTBLS is sourced from scientific articles hosted on arXiv  [^2] https://arxiv.org, an open-access repository of academic preprints.
Beyond factoid question-answering, iTBLS encompasses tasks such as comparison, determining absolute and relative positions, and mathematical reasoning. Previous research primarily examined procedural command generation for spreadsheets or the alignment of tabular data through LLMs. iTBLS integrates these functionalities into a unified task, enabling the manipulation of existing tables through natural-language commands. On tabular generation, while prior work addressed the summarization of natural language paragraphs in a tabular format, iTBLS focuses on generating row or column data conversationally.
In addition to building iTBLS, this paper develops a novel approach to address tabular operations by reformulating the task as conditional question answering. Furthermore, the question-answering-based reformulation is applied to other datasets introduced in prior work Wu et al. (2022) and results in better performance in terms of both table-cell accuracy and BERTScore.
@@
  • Creating iTBLS, a dataset of tabular interactions unifying interpretation, modification, and generation.
@@
  • Broadening the scope of interactions to include mathematical reasoning, natural language manipulation, and natural language expansion.
  • Introducing a novel approach for table generation tasks through a two-stage reformulation that first identifies the cells to be manipulated and generates a question based on the requested operation, then answers those questions using the user request and the input table as evidence.
  • Demonstrating up to 13% improvement in table-cell accuracy and up to 16% improvement in BERTScore using the novel approach on the text-to-table task introduced by prior work.
@@ 2.1 Tabular Interpretation
Tabular interpretation involves a dialogue turn focused on extracting information from a specific cell in a table, such as identifying a cell satisfying certain criteria. Prior research on tabular interpretation focused on grounded question-answering. An important challenge in the collection of such datasets is the availability of large-scale tabular data. Consequently, many tabular datasets are constructed from online resources such as Wikipedia including WikiTableQuestions Pasupat and Liang (2015), ManyModalQA Hannan et al. (2020), TaBERT Yin et al. (2020), NQ-Tables Herzig et al. (2021), FEVEROUS Aly et al. (2021), FeTaQA Nan et al. (2022), HybriDialogue Nakamura et al. (2022b), and HiTab Cheng et al. (2022). Other tabular datasets are constructed from financial reports including TAT-QA Zhu et al. (2021), FinQA Chen et al. (2021), MultiHiertt Zhao et al. (2022), or scientific reviews Sundar et al. (2024).
Proposed approaches to address the tabular interpretation task include architectures based off of the Transformer encoder Yin et al. (2020); Herzig et al. (2020); Chen et al. (2019b); Eisenschlos et al. (2020); Liu et al. (2021); Gu et al. (2022); Yang et al. (2022), decoder Gong et al. (2020); Akhtar et al. (2023); Zha et al. (2023); Jiang et al. (2023); Zhang et al. (2023); Sui et al. (2024); Cremaschi et al. (2025), or both (encoder-decoder) Nakamura et al. (2022b); Deng et al. (2022); Sundar and Heck (2023).
@@ 2.2 Tabular Modification
Tabular modification concerns the manipulation of the content within an existing table without altering the overall structure of rows and columns. Early work on tabular modification explored the generation of procedural commands for spreadsheets using synthesis algorithms Singh and Gulwani (2012); Shigarov et al. (2019). Tools utilizing programming-by-example to parse user intents into executable commands have also been explored Scaffidi et al. (2009); Kandel et al. (2011); Jin et al. (2017); Petricek et al. (2023); Chen et al. (2023); Xing et al. (2024). More recent work has shifted focus towards leveraging LLMs to synthesize commands for tools Huang et al. (2024), reformat tabular information Dargahi Nobari and Rafiei (2024), and execute programming commands Liu et al. (2024).
@@ 2.3 Tabular Generation
Tabular generation focuses on expanding an existing table by adding a new row or column. Research on tabular generation initially employed discriminative techniques, such as tree-based methods for generating tables of contents Branavan et al. (2007) and SVMs to classify text across various labels Aramaki et al. (2009). Recent approaches have shifted towards neural techniques including Generative Adversarial Networks (GANs) Xu and Veeramachaneni (2018); Park et al. (2018); Chen et al. (2019a); Zhao et al. (2021), Autoencoders Li et al. (2019); Darabi and Elor (2021), Diffusion models Kotelnikov et al. (2023), and LLMs Borisov et al. (2023); Solatorio and Dupriez (2023); Gulati and Roysdon (2023); Zhao et al. (2023); Seedat et al. (2024); Deng et al. (2024).
A similar line of research also explores the generation of tabular data from associated textual information. Wu et al. (2022) introduced four datasets and proposed a modification to the Transformer’s attention mechanism to summarize textual information in a tabular format by inverting datasets created for the dual task of converting tables to text, (as opposed to new conversational evidence). Other approaches to summarize textual information in a tabular format include the addition of learnable bias parameters Pietruszka et al. (2022) and structure-aware instruction-tuning Tang et al. (2023).
@@ 3 The iTBLS Dataset
The Interactive Tables (iTBLS) dataset features conversational interactions situated in tabular data, covering the three distinct types of interactions described in Section 2: interpretation, modification, and generation. Each example type is exemplified in Figure 1 and described below. In addition, since the mode of interaction is not known a priori, any proposed approach using iTBLS must effectively identify the interaction type, either explicitly or implicitly. In the following sections, we provide a detailed description of each type of interaction and outline the dataset collection process.
@@ 3.1 Tasks
Tabular Interpretation: In iTBLS, interpretive interactions are structured as question-answer pairs, where the goal is to identify the cell referred to by the question. The references could be absolute (referring to a specific row or column), or relative (referring to one cell in the context of another). Appendix LABEL:sec:dataset_examples_itbls details absolute and relative references in iTBLS.
Tabular modification: We conceptualize modification in iTBLS as a series of cell swaps, positing that any content rearrangement can ultimately be reduced to such exchanges. This approach allows for both explicit references, where specific row and column numbers are cited, and implicit references, which rely on the content or relative positions of cells. Table LABEL:tab:modif_example in Appendix LABEL:sec:dataset_examples_itbls showcases examples from iTBLS. As observed, there is a mix of explicit and implicit references to the specific contents to be manipulated.
Tabular generation: In iTBLS, table generation is guided by new natural language evidence. This evidence clarifies appending a row or column, defines the suitable header, and supplies the data entries for the new row (or column) relative to existing columns (or rows). This process ensures that the added elements are contextually relevant and accurately integrated into the table. Table LABEL:tab:generat_example in Appendix LABEL:sec:dataset_examples_itbls provides examples of such interactions, demonstrating how users can request the incorporation of new row and column data into an established table framework.
@@ 3.2 Dataset Collection
For our dataset creation, we employ two distinct sources for generating dialogue turns based on the type of interaction and the specific table component involved. For tasks related to tabular interpretation and modification, we engage crowd-workers from Amazon Mechanical Turk (AMT). These workers are tasked with formulating questions or commands that pertain to the pre-identified cell(s) designated for each interaction. We recruit workers from Australia, Canada, Ireland, New Zealand, the United Kingdom, and the USA. Each crowdworker is compensated at a rate of \$0.15 per Human Intelligence Task (HIT), with the average completion time for each HIT being approximately 40 seconds. Detailed information on the AMT interface used for these tasks is included in Appendix LABEL:sec:app_mturk.
For generation, GPT-4 is prompted to write a dialogue turn summarizing a row or column of the table. The prompt is as follows:
@@
Each sample in the dataset contains the source arXiv ID, the table that the conversation is situated in, the index of that table within the paper (e.g. Table X), the utterance describing the interaction, the ground truth cell(s) involved in the interaction, and finally the expected output. Statistics of the datasets are provided in Table 1.
Statistic Interpret Modify Generate
# Samples 4168 4168 4168
@@
Words 10.6 13.4 31.6
Tokens 14.3 18.3 59.1
# Per table
Cells 28.1 28.1 25.31
(Cols/Rows) 5.0/5.5 5.0/5.5 4.8/5.3
Table 1: Statistics of the iTBLS dataset
4 Methods
4.1 Table operations through conditional question answering
We also present a novel approach that reformulates operations on tables as question answering. A primary challenge in tabular operations using LLMs lies in ensuring the syntactic validity of the produced tables. Every row and column in a table must contain the same number of cells, with row and column headers delineating relationships between cells. Failing to adhere to this constraint invalidates the structure of the table and the information presented. Prior work addresses this constraint by including additional parameters like row and column relation embeddings Wu et al. (2022) or positional bias Pietruszka et al. (2022) to get the model to attend to header cells while generating content. However, this results in highly specialized architectures for a singular task. Breaking the task down into question-answering results in a more interpretable framework while ensuring validity of the generated tables.
The first step identifies the mode of interaction and the cell(s) the user is referring to, which is used to formulate a question. The second step converts the table into a pandas dataframe, parses the table and the question generated from the previous step to obtain a pandas command corresponding to the task, and executes the command on the dataframe to generate the final table. Generating a valid command ensures that the final table is syntactically valid as well (that is, the number of columns across all rows is consistent).
For the interpret task, the question-answer reformulation is trivial, since all interpretive queries and associated responses are naturally question-answer pairs. For the modify task, the question is of the form To which cells is the user referring?. A language model is then fine-tuned to generate a response containing the cells (indexed by row and column). Then, the LLM response is reformatted into an appropriate pandas command. Finally, for the generate task, the question-answering is more nuanced. First, the user request is parsed to identify whether a row or a column is to be appended. The header of the corresponding row is then extracted from the user request. Using the extracted header and the other header cells of the table, questions are generated for each of the empty cells to be filled in the form What is the row value for column?. The user request is parsed to obtain the answers to these generated questions, forming the corresponding row or column to be appended.
Figure 2: Overview of the novel question-answering reformulation to perform table operations